# Calibration Attacks: A Comprehensive Study of Adversarial Attacks on Model Confidence

### v2

@@ 4.4 Insights on Key Aspects of Attacks
We analyze key aspects of calibration attacks: attack directions, noise bounds, and attack iterations under the $l_{\infty}$-based attacks. More details can be found in Appendix G, including attack effectiveness on data with varying imbalance ratios.
@@
Perturbation Noise Levels. We study the effect of perturbation noise levels and how low the value could be in order to construct notable harm. As shown in the left three subfigures in Figure 2, under different setups, the attacks are highly successful even at a low level of noise (e.g., $\epsilon=0.05$ or even lower). We can also see the rise in ECE is sharp when $\epsilon$ increases, and it plateaus quickly.
Iterations. The numbers of iterations can help measure the efficiency and cost of attacks (in terms of both time and computing expenditure). The last subfigure in Figure 2 shows ECE vs. iteration numbers when applying MMA to the victim models on CIFAR-100. In Appendix G.3, we provide more detailed comparison, showing that the calibration attacks consistently produced a higher degree of miscalibration compared to the original unaltered SA. ECE begins to saturate at 500 iterations, but even at 100 iterations the victim models become heavily miscalibrated.
Table 3: Comparison between the efficiency of the underconfidence (shortened as und.) and overconfidence (shortened as ovr.) attacks.
@@ 4.5 Qualitative Analysis
In Figure 3 we show the coarse localization maps produced with GradCAM (Selvaraju et al., 2017), which highlights the most important regions that a model relies on to make prediction using the gradients from different layers of a network. We apply GradCAM to our ResNet-50 models that are fine-tuned and tested on Caltech-101, based on the standard attack settings discussed in Section C. We choose images where the attacks led to large change in predicted confidence (at least 10%). Figure 3 shows several representative images. We can see that the coarse localization maps have minimal to no noticeable changes after the adversarial images are produced, especially in the case of the overconfidence attacked images. This analysis shows that it could be difficult to identify the attacks based on gradient visualization methods.
Figure 3: The GradCAM visualizations shows the image regions most responsible for the decisions of ResNet-50 before (top row) and after (bottom row) attacks. The left three images are under underconfidence attacks and the right three the overconfidence attack.
@@
Appendix G.4 provides detailed t-SNE visualization on the effect of different forms of calibration attacks. The visualization shows that overconfidence attack causes the representations for different classes to be split apart as much as possible, while the underconfidence attack causes a more jumbled representation with most data points falling closely to the decision boundary. The visualization shows that the attacks achieve their intended goals.
4.6 Confidence Certification.
@@
We follow the method in Kumar et al. (2020) to compare the lower bound of expected confidence of a smoothed classifier between base images and their underconfidence and overconfidence attacked counterparts. We find that the lower bound is not significantly different between adversarially attacked images and their originals, though the accuracy of the smoothed classifier is affected. This means that methods for determining certified lower bounds are not strongly affected by calibration attacked samples (similar to conventional adversarial samples), but limitations in efficiency and accuracy of these methods means that establishing defence method against such attacks is still paramount. A detailed discussion is in Appendix G.7.
@@ 5 Defending Against Calibration Attacks
We compare a wide range of recalibration and defence methods under the setup of the MMA attacks, which, as shown above, are among the most effective calibration attacks.
Specifically, for post-calibration methods, we include Temperature Scaling (TS) (Guo et al., 2017), Multi-domain Temperature Scaling (MD-TS) (Yu et al., 2022), and calibration with splines (Splines) (Gupta et al., 2021). For training-based regularization methods we include two effective models, DCA (Liang et al., 2020) and SAM (Foret et al., 2021). Regarding adversarial defence methods, we test the top-3 SOTA models under the $l_{\infty}$ attack for CIFAR-100, using WideResNet on the RobustBench leaderboard (Croce et al., 2021), which are only available for CIFAR-10 and CIFAR-100, hence we run over CIFAR-100 to compare with our previous baselines. We further include a recent post-process defence called Adversarial Attack Against Attackers (AAA) (Chen et al., 2022) in addition to the the most common defences in the form of a PGD-based adversarial training (AT) (Madry et al., 2018). We also included the two defences proposed for calibrations attacks: CAAT and CS, which are introduced in Section 3.3. (The experiment setup is described in Appendix C, and the detailed description of these baselines are in Appendix D.2).
Results and Analyses. Table 4 shows the experiment results of query efficiency, accuracy, and the ECE and KS errors, before (PrECE and PrKS) and after the attacks (PsECE and PrKS), using different recalibration and defence models. The best performances of post-attack ECE (PsECE) and KS (PsKS) are marked in bold, and the second best are marked with underlines.
We can see that CS is overall the strongest methods at maintaining low calibration errors on post-attack datapoints. It showed the best post-attack calibration performance in five out of six setups on PsECE, and ranked among top 2 in all the other setups. The other top calibration performances are distributed among AAA, Spline, and AT. Knowing the property of calibration attacks and organizing defences accordingly is helpful to ensure a better defending result.
@@ 6 Conclusions
We highlight and perform a comprehensive and dedicated study on calibration attacks, which aim to trap victim models into being heavily miscalibrated, hence endangering the trustworthiness of the models and any follow-up decision-making processes based on confidences. We propose four typical forms of calibration attacks and demonstrate their severity from different perspectives. We also show calibration attacks are difficult to detect compared to standard attacks. Investigation is then conducted to study the effectiveness of a wide range of adversarial defences and calibration methods, including the defences that are specifically designed for calibration attacks. From the ECE and KS scores, we can see that there are still limitations on these recalibration and defences in handling calibration attacks. We hope this paper helps attract more attention to the attacks against confidence and hence mitigate their potential harm. We provides detailed analyses to help understand the characteristics of the attacks for future work.
@@ Appendix A Detailed Summary of Related Work
In the field of calibration, a great deal of current research is devoted to the creation of new calibration methods that can be applied to create better calibrated models while possessing as minimum overhead in applying them as possible. Methods are generally divided into two types. Post-calibration methods can be applied directly to the predictions of fully trained models at test time, and methods of this class include temperature scaling (Guo et al., 2017). More traditional methods of this type include Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002), and histogram binning (Zadrozny & Elkan, 2001). All of these three methods are originally formulated for binary classification settings, and work by creating a function that maps predicted probabilities based on their values more in tune with the model’s level of performance. Although they are easy to apply, they often come with the limitation of needing a large degree of validation data to tune, especially with isotonic regression, and performance can struggle when applied to more out of distribution data.
@@
With respect to adversarial attacks, attacks in this field are wide ranging. Well known white-box attacks include the basic Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015). This method works by finding adjustments to the input data that maximizes the loss function, and uses the backprogogated gradients to produce the adversarial examples. Projected gradient descent (PGD) (Madry et al., 2018) is popular iterative-based method that similarly uses gradient information, and has been shown to be a universal first-order adversary, and thus is the strongest form of attacks making using of gradient and loss information. In the black-box space of attacks, ZOO (Chen et al., 2017) is an example of a popular score-based attack that uses zeroth order stochastic coordinate descent to attack the model, and avoids training a substitute model. The authors make use of attack-space dimension reduction, hierarchical attacks and importance sampling to make the attack more query efficient, which is required as black-box attacks generally need a lot of queries to run compared to white-box methods.
@@
Appendix B Details for Maximum Miscalibration Attacks
@@
Appendix C Details of Experimental Setup
Metrics. To assess the degree of calibration error caused by each attack, we use two metrics, the popular binning-based Expect Calibration Error (ECE) (Pakdaman Naeini et al., 2015), and KS errors (Gupta et al., 2021), which are formulated in detail in Section E.
@@ Appendix C Details of Experimental Setup
Models. ResNet-50 (He et al., 2016) is primary model we train and test on due to it being a standard model for image classification. Non-convolutional attention-based networks have recently attained great results on image classification tasks, so we also experiment with the popular Vision Transformer (ViT) architecture (Dosovitskiy et al., 2021). Both of these models are the versions with weights pretrained on ImageNet (Deng et al., 2009). We use the VIT_B_16 variant of ViT, and the pretraining dataset used for each model is ImageNet_1K for ResNet and ImageNet_21K for ViT, and are fine-tuned on the target datasets. Pretrained models are advantageous to study given they can increase performance over training from randomly initialized weights and is a more practical use-case. The specific details behind our training procedures and our various model hyperparameters can be seen in Section D.
Attack Settings. Regarding the SA version of the attacks, for the $l_{\infty}$ and $l_{2}$ norm attacks we use the default SA settings for $\epsilon$ and $p$, which are $\epsilon=0.05$ and $p=0.05$ for $l_{\infty}$ and $\epsilon=5.0$ and $p=0.1$ for $l_{2}$. For our primary results we run the attacks on a representative 500 test cases from the test set of each dataset. Each attack is ran for 1000 iterations, far less than the default 10,000 in Andriushchenko et al. (2020), but since there is no need to change the label, less iterations are required, bolstering the use-case and threat for this form of attacking.
@@
Appendix D Specific Training Details
D.1 General Settings

### v3

@@ 4.4 Insights on Key Aspects of Attacks
We analyze key aspects of calibration attacks: attack directions, noise bounds, and attack iterations under the $l_{\infty}$-based attacks. More details can be found in Appendix H, including attack effectiveness on data with varying imbalance ratios.
@@
Perturbation Noise Levels. We study the effect of perturbation noise levels and how low the value could be in order to construct notable harm. As shown in the left three sub-figures in Figure 2, under different setups, the attacks are highly successful even at a low level of noise (e.g., $\epsilon=0.05$ or even lower). We can also see the rise in ECE is sharp when $\epsilon$ increases, and it plateaus quickly.
Iterations. The numbers of iterations can help measure the efficiency and cost of attacks (in terms of both time and computing expenditure). The last sub-figure in Figure 2 shows ECE vs. iteration numbers when applying MMA to the victim models on CIFAR-100. In Appendix H.3, we provide more detailed comparison, showing that the calibration attacks consistently produced a higher degree of miscalibration compared to the original unaltered SA. ECE begins to saturate at 500 iterations, but even at 100 iterations the victim models become heavily miscalibrated.
Figure 3: The GradCAM visualizations shows the image regions most responsible for the decisions of ResNet-50 before (top row) and after (bottom row) attacks. The left three images are under UCA and the right three the OCA.
Table 3: Comparison between the efficiency of the UCA and OCA.
@@ 4.5 Qualitative Analysis
In Figure 3 we show the coarse localization maps produced with GradCAM (Selvaraju et al., 2017), which highlights the most important regions that a model relies on to make prediction using the gradients from different layers of a network. We apply GradCAM to our ResNet-50 models that are fine-tuned and tested on Caltech-101, based on the standard attack settings discussed in Section D. We choose images where the attacks led to large change in predicted confidence (at least 10%). Figure 3 shows several representative images. We can see that the coarse localization maps have minimal to no noticeable changes after the adversarial images are produced, especially in the case of the overconfidence attacked images. This analysis shows that it could be difficult to identify the attacks based on gradient visualization methods.
@@
Appendix H.4 provides detailed t-SNE visualization on the effect of different forms of calibration attacks. The visualization shows that the overconfidence attack causes the representations for different classes to be split apart as much as possible, while the underconfidence attack causes a more jumbled representation with most data points falling closely to the decision boundary. The visualization shows that the attacks achieve their intended goals.
4.6 Confidence Certification
@@
We follow the method in Kumar et al. (2020) to compare the lower bound of expected confidence of a smoothed classifier between base images and their under- and overconfidence attacked counterparts. We find that the lower bound is not significantly different between adversarially attacked images and their originals, though the accuracy of the smoothed classifier is affected. This means that methods for determining certified lower bounds are not strongly affected by calibration attacked samples (similar to conventional adversarial samples), but limitations in efficiency and accuracy of these methods means that establishing defence method against such attacks is still paramount. A detailed discussion is in Appendix H.7.
4.7 Attack Transferability
Table 4: Attack transfer ViT to ResNet and to ensembled victim models (with ViT as target) on CIFAR-100.
ECE KS Avg.-Conf.
Pre-atk .$051_{\pm.005}.$$035_{\pm.006}.$$916_{\pm.006}.$
UCA SQ VIT → ResNet .$104_{\pm.003}.$$101_{\pm.007}.$$780_{\pm.012}.$
UCA PGD VIT→ ResNet .$196_{\pm.020}.$$194_{\pm.021}.$$729_{\pm.011}.$
Pre-atk .$028_{\pm.005}.$$015_{\pm.008}.$$920_{\pm.002}.$
UCA SQ Ensemble .$091_{\pm.010}.$$086_{\pm.008}.$$708_{\pm.012}.$
UCA PGD Ensemble .$048_{\pm.011}.$$014_{\pm.006}.$$637_{\pm.030}.$
Determining the transferability of calibration attacks across different model architectures is of interest since although one can effectively optimize attacks against a single model, using an ensemble model could potentially be a form of defence if transferability is limited. We investigate how our attacks transfer to different victim models. Specifically in our experiments, the attacks are constructed on CIFAR-100 images with regular settings based on ViT but used on (i) ResNet and (ii) ensemble victim models combining prediction of both ResNet-50 and ViT (Table 4). We focus on the transfer of the UCA attack, given that OCA has little room to skew confidence due to the already high accuracy of victim models. The attacks show transferability, although in general they become less effective when applied on different architectures. This is expected because the attacks are constructed without considering the target victim models’ properties. Most notably, the simple ensemble of two typical victim models (ViT and ResNet) is still greatly subject to the attacks (the output distributions of ResNet and ViT are averaged to get the ensemble), so using ensembles as a defence does not fully protect against calibration attacks. To further explore transferability, attack methods need to be explicitly redesigned (Xie et al., 2018; Wang et al., 2021), which could be further studied in the future.
4.8 Influence of Model Scale on Attack Effectiveness
Table 5: Attack effectiveness of MMA (black-box setup) across different model sizes on CIFAR-100.
Avg #q Med. #q ECE KS Avg.Conf.
ResNet-18 (Acc: .$854_{\pm.004}.$)
Pre - - .$037_{\pm.003}.$$016_{\pm.005}.$$870_{\pm.001}.$
Post $52.9_{\pm 1.4}$ $36.0_{\pm 0.0}$$531_{\pm.009}.$$471_{\pm.004}.$$446_{\pm.003}.$
ResNet-50 (Acc: .$881_{\pm.002}.$)
Pre - - .$052_{\pm.006}.$$035_{\pm.006}.$$916_{\pm.006}.$
Post $72.9_{\pm 2.8}$ $41.5_{\pm 2.8}$$606_{\pm.002}.$$497_{\pm.002}.$$502_{\pm.002}.$
ResNet-152 (Acc: .$883_{\pm.002}.$)
Pre - - .$051_{\pm.005}.$$046_{\pm.004}.$$929_{\pm.005}.$
Post $80.5_{\pm 3.8}$ $47.0_{\pm 3.5}$$539_{\pm.002}.$$475_{\pm.002}.$$472_{\pm.008}.$
ViT-base (Acc: .$935_{\pm.002}.$)
Pre - - .$064_{\pm.006}.$$054_{\pm.005}.$$882_{\pm.004}.$
Post $104.8_{\pm 7.5}$ $62.7_{\pm 4.7}$$616_{\pm.003}.$$564_{\pm.000}.$$431_{\pm.001}.$
ViT-large (Acc: .$936_{\pm 0.001}.$)
Pre - - .$037_{\pm.009}.$$022_{\pm.012}.$$921_{\pm.021}.$
Post $122.1_{\pm 2.0}$ $76.3_{\pm 3.2}$$531_{\pm.012}.$$508_{\pm.020}.$$450_{\pm.024}.$
To help understand the influence of varying model sizes on attack performance, we perform a study on models of different sizes over different architectures across four attack variants. Specifically, for ResNet, we compare ResNet-18, ResNet-50, and ResNet-152. For ViT, we included Base (vit-base-patch16-224-in21k) and Large models (vit-large-patch16-224-in21k). The model training closely follows the settings detailed in Appendix D.1. We fine-tune the models for each variant and average the results. We ran attacks with the same settings used to produce Tables 1 and 3.
In Table 5 we present the results of the strongest variant of calibration attack (i.e., MMA). It shows that the proposed calibration attacks are effective among different model sizes. The model sizes do not have a major impact on the susceptibility to calibration attacks, as the overall ECE and KS scores do not follow a clear pattern along with varying model sizes. However, when models scale up, the numbers of queries increase in general, suggesting that attacking complex models requires more rounds of attacks in most cases. In Appendix H.11 we present detailed results in the black-box and white-box setups across three architectures under the four types of attacks, which agree with our aforementioned observations.
@@ 5 Defending Against Calibration Attacks
We compare a wide range of recalibration and defence methods under the setup of the maximum miscalibration attacks, which, as shown above, are among the most effective calibration attacks.
Specifically, for post-calibration methods, we include Temperature Scaling (TS) (Guo et al., 2017), Multi-domain Temperature Scaling (MD-TS) (Yu et al., 2022), and calibration with splines (Splines) (Gupta et al., 2021). For training-based regularization methods we include two effective models, DCA (Liang et al., 2020) and SAM (Foret et al., 2021). Regarding adversarial defence methods, we test the top-3 SOTA models under the $l_{\infty}$ attack for CIFAR-100, using WideResNet on the RobustBench leaderboard (Croce et al., 2021), which are only available for CIFAR-10 and CIFAR-100, hence we run over CIFAR-100 to compare with our previous baselines. We further include a recent post-process defence called Adversarial Attack Against Attackers (AAA) (Chen et al., 2022) in addition to the the most common defences in the form of PGD-based adversarial training (AT) (Madry et al., 2018). We also included the two defences proposed for calibrations attacks: CAAT and CS, which are introduced in Section 3.3. (The experiment setup is described in Appendix D, and the detailed description of these baselines are in Appendix E.2).
Results and Analyses. Table 6 shows the experiment results of query efficiency, accuracy, and the ECE and KS errors, before (PrECE and PrKS) and after the attacks (PsECE and PrKS), using different recalibration and defence models. The methods with the best performances in terms of post-attack ECE (PsECE) and KS (PsKS) are marked in bold, and the second best are underlined.
We can see that CS is overall the strongest methods at maintaining low calibration errors on post-attack datapoints. It showed the best post-attack calibration performance in five out of six setups on PsECE, and ranked among top 2 in all the other setups. The other top calibration performances are distributed among AAA, Splines, and AT. Knowing the property of calibration attacks and organizing defences accordingly is helpful to ensure better defence results.
@@ 6 Conclusions
We highlight and perform a comprehensive and dedicated study on calibration attacks, which aim to trap victim models into being heavily miscalibrated, hence endangering the trustworthiness of the models and any follow-up decision-making processes based on confidences. We propose four forms of calibration attacks and demonstrate their severity from different perspectives. We also show calibration attacks are difficult to detect compared to standard attacks. An investigation was conducted to study the effectiveness of a wide range of adversarial defences and calibration methods, including the defences that are specifically designed for calibration attacks. From the ECE and KS scores, we can see that there are still limitations on these recalibration and defences in handling calibration attacks. We hope this paper helps attract more attention to the attacks against confidence and hence mitigate their potential harm. We provides detailed analyses to help understand the characteristics of the attacks for future work.
@@ Appendix A Detailed Summary of Related Work
In the field of calibration, a great deal of current research is devoted to the creation of new calibration methods that can be applied to create better calibrated models while possessing as minimum overhead in applying them as possible. Methods are generally divided into either post-calibration or training-based. Post-calibration methods can be applied directly to the predictions of fully trained models at test time, and methods of this class include temperature scaling (Guo et al., 2017). More traditional methods of this type include Platt scaling (Platt, 1999), isotonic regression (Zadrozny & Elkan, 2002), and histogram binning (Zadrozny & Elkan, 2001). All of these three methods are originally formulated for binary classification settings, and work by creating a function that maps predicted probabilities based on their values more in tune with the model’s level of performance. Although they are easy to apply, they often come with the limitation of needing a large degree of validation data to tune, especially with isotonic regression, and performance can struggle when applied to more out of distribution data.
@@
With respect to adversarial attacks, attacks in this field are wide ranging. Well known white-box attacks include the basic Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015). This method works by finding adjustments to the input data that maximizes the loss function, and uses the backpropagated gradients to produce the adversarial examples. Projected gradient descent (PGD) (Madry et al., 2018) is popular iterative-based method that similarly uses gradient information, and has been shown to be a universal first-order adversary, and thus is the strongest form of attacks making using of gradient and loss information. In the black-box space of attacks, ZOO (Chen et al., 2017) is an example of a popular score-based attack that uses zeroth order stochastic coordinate descent to attack the model, and avoids training a substitute model. The authors make use of attack-space dimension reduction, hierarchical attacks and importance sampling to make the attack more query efficient, which is required as black-box attacks generally need a lot of queries to run compared to white-box methods.
@@
Appendix B Miscalibration Effects on Downstream Tasks
In terms of specific real-world applications, skewed confidence scores or uncertainties have been shown to have dramatic impacts on downstream tasks and human decision-making, which many previous works have highlighted. Previous studies such as those of Penso et al. (2024); Jiang et al. (2012); Kompa et al. (2021) have shown the effect of miscalibration on down-stream authority mechanisms. Penso et al. (2024) discuss how medical imaging, if skipping expert review for the confident but incorrect model predictions, can have disastrous consequences, stressing the importance of being well-calibrated on incorrect predictions. Kompa et al. (2021) advocate for medical AI models being able to abstain from providing a diagnosis when there is a significant uncertainty for a patient, and to defer to additional human expertise to make a more informed diagnosis. This relies on a model that can give well-calibrated estimates.
Moreover, prior works have examined the role of skewed confidence affecting the level of trust humans have in a system, which influences applications where humans and AI decision-making occurs symbiotically Zhang et al. (2020); Rechkemmer & Yin (2022). Antifakos et al. (2005), through a study on context-aware mobile phones, show how providing explicit confidence improves user trust in a system. Zhang et al. (2020) demonstrate when a model produces high confidence scores, users are more prone to rely on the AI for decision-making. These factors show why miscalibrated outputs can have great influence on downstream processing and authority mechanisms.
As more specific examples in the healthcare domain, chest abnormality detection Dyer et al. (2021) tag radiographs with high-confidence normality to help radiologists make decisions. In child maltreatment hotlines, screening models De-Arteaga et al. (2020) are deployed where confidence is provided to suggest the likelihood that a child on the referral will experience adverse welfare-related outcomes. Call workers use confidence, along with other information, to make their decisions. The study discusses how the tool could be “glitching” out by providing incorrect confidence, compromising operability. In the legal domain, LegalLens Bernsohn et al. (2024) detects legal violations. Entities with high confidence are further processed with downstream models. In enterprise security, anomalous behaviour detectors Shashanka et al. (2016) use confidence to send alerts about likely malicious activity.
Dhuliawala et al. (2023) conduct user studies that show that confidently incorrect predictions heavily undermine user trust in human-AI collaboration settings, requiring only a few incorrect instances with inaccurate confidence estimates to damage user trust and performance, creating long-term effects with recovery being very difficult once humans have been exposed to miscalibrated confidence estimates. They observe that both confidently incorrect stimuli and unconfidently correct stimuli reduce trust, with the former being more significant, which justifies our formulation of MMA as having a high potential for creating mistrust by optimizing the creation of both of these stimuli.
Appendix C Details for Maximum Miscalibration Attacks
@@
Appendix D Details of Experimental Setup
Metrics. To assess the degree of calibration error caused by each attack, we use two metrics, the popular binning-based Expected Calibration Error (ECE) (Pakdaman Naeini et al., 2015), and KS error (Gupta et al., 2021), which are formulated in detail in Section F.
@@ Appendix D Details of Experimental Setup
Models. ResNet-50 (He et al., 2016) is the primary model we train and test on due to it being a standard model for image classification. Non-convolutional attention-based networks have recently attained great results on image classification tasks, so we also experiment with the popular Vision Transformer (ViT) architecture (Dosovitskiy et al., 2021). Both of these models are the versions with weights pretrained on ImageNet (Deng et al., 2009). We use the VIT_B_16 variant of ViT, and the pretraining dataset used for each model is ImageNet_1K for ResNet and ImageNet_21K for ViT, and are fine-tuned on the target datasets. Pretrained models are advantageous to study given they can increase performance over training from randomly initialized weights and present a more practical use-case. The specific details behind our training procedures and our various model hyperparameters can be seen in Section E.
Attack Settings. Regarding the SA version of the attacks, for the $l_{\infty}$ and $l_{2}$ norm attacks we use the default SA settings for $\epsilon$ and $p$, which are $\epsilon=0.05$ and $p=0.05$ for $l_{\infty}$ and $\epsilon=5.0$ and $p=0.1$ for $l_{2}$. For our primary results we run the attacks on a representative 500 test cases from the test set of each dataset. Each attack is ran for 1000 iterations, far less than the default 10,000 in Andriushchenko et al. (2020), but since there is no need to change the label, less iterations are required, bolstering the use-case and threat for this form of attack.
@@
Appendix E Specific Training Details
E.1 General Settings